# optimizn: a Python Library for Developing Customized Optimization Algorithms

### v1

@@ 6 Applicability to a Niche $NP$-Hard Problem
  • An undirected graph $G=(V=V_{1}\cup\dots\cup V_{d},E)$ where the vertices in $V$ are split into $d$ groups, each representing a testing dimension (e.g. hardware model, virtual machine type). Each vertex is referred to as a “dimension value”. The edges in $E$ represent compatibility relationships between the vertices in $V$. Dimension values in the same dimension are considered incompatible $\forall k\in\{1,\dots,d\},\forall v_{i},v_{j}\in V_{k},(v_{i},v_{j})\not\in E$
  • The number of testing configurations $n\in\mathbb{Z}^{+}$.
  • Objective function $O(S)$, that when given a collection of testing configurations (“schedule”) $S$, returns a non-negative number that quantifies how close the distribution of dimension values, pairs of dimension values, or testing configurations in $S$ is to the desired/target distribution ($O(S)=0$ indicates a perfect match).
  • Scope $C=(I=I_{1}\cup\dots\cup I_{d},X=X_{1}\cup\dots\cup X_{d})$, which consists of collections of values to include ($I$) and exclude ($X$) for each dimension.
@@
  • Each testing configuration has $d$ dimension values, $\forall s\in S,|s|=d$

### v2

@@ 6 Applicability to a Niche $NP$-Hard Problem
  1. 1.
    An undirected graph $G=(V=V_{1}\cup\dots\cup V_{d},E)$ where the vertices in $V$ are split into $d\geq 2$ groups, each group representing a testing dimension (e.g. HW model, VM type) and each vertex representing a dimension value. The edges in $E$ represent compatibility relationships between the vertices in $V$. Dimension values in the same dimension are considered incompatible $\forall k\in\{1,\dots,d\},\forall v_{i},v_{j}\in V_{k},(v_{i},v_{j})\not\in E$
  2. 2.
    The number of testing configurations $n\in\mathbb{Z}^{+}$
  3. 3.
    Objective function $O(S)$, that when given a collection of testing configurations (“schedule”) $S$, returns a non-negative number that quantifies how close the target distribution of dimension values is to the distribution of dimension values in $S$ (0 indicates a perfect match).
  4. 4.
    Scope $C=(I=I_{1}\cup\dots\cup I_{d},X=X_{1}\cup\dots\cup X_{d})$, $\forall i,I_{i}\cap X_{i}=\emptyset$, which consists of collections of dimension values (for each dimension) that must be included in ($I$) and excluded from ($X$) the output.
@@
  2. 2.
    Each testing configuration has $d$ dimension values $\forall s\in S,|s|=d$, exactly one dimension value from each dimension $\forall s\in S,\forall i\in\{1,\dots,d\},|s\cap V_{i}|=1$
  3. 3.